# Localized excitation on the Jacobi elliptic periodic background for the (n+1)-dimensional generalized Kadomtsev-Petviashvili equation

### v1

@@ 6 Degenerate solutions
In this section, we discuss the degenerate cases of the periodic wave background, that is the Jacobi elliptic function with a modulus $k$ approaching either 00 or $1$, respectively. To ensure that the solutions are real after degeneracy, we consider the breather waves (53) and (65), and $\lambda\in[-k^{2}-\delta\omega_{2},+\infty)$ in both cases.
@@
×$\quad\quad~{}\times\cosh(-\bar{\kappa}_{1}+\bar{\kappa}_{2}))^{2},$
@@
In this case, $E(k)/K(k)=0$, $F(\phi_{\gamma},k)\to\infty$ can be derived. Based on the translation invariance of the (n+1)-dimensions gKP equation (1), the half-period translation of (53) with the transformation $\eta=\eta-K(k)$ is defined to ensure the standardization of degradation solution. On account of the fact that $\gamma=F(\phi_{\gamma},k)$, for each $\lambda\in[-1-\delta\omega_{2},+\infty)$, $\chi_{1}$ can be defined according to its property [66] as
@@
At the same time, employing the Poisson summation equation [67], if $k=1$,
@@
×$\quad\quad~{}\times\mathrm{exp}(-\bar{\kappa}_{1}))^{2},$
@@
×$\quad\ \ \left.\times\mathrm{cosh(\eta-\chi_{2})}\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})+\mathrm{cosh(\eta+\chi_{1})}\right.$
×$\quad\ \ \left.\times\mathrm{cosh(\eta+\chi_{2})}\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2}))+(\bar{\mu}_{1}+\bar{\mu}_{2})\right.$
×$\quad\ \ \left.\times(\mathrm{cosh(\eta+\chi_{1})}\mathrm{cosh(\eta-\chi_{2})}\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right.$
@@
$\bar{F}_{2}=\frac{-2k^{\prime}\ln k^{\prime}}{\pi}\left((-\bar{\mu}_{1}+\bar{\mu}_{2})\left(2\sinh(\chi_{1}-\eta)\right.\right.$
×$\quad\ \ \left.\left.\times\sinh(\chi_{2}-\eta)-(\bar{\mu}_{1}+\bar{\mu}_{2})\left(\sinh(\chi_{1}-\eta)\right.\right.\right.$
×$\quad\ \ \left.\left.\left.\times\cosh(\chi_{2}-\eta)+\sinh(\chi_{2}-\eta)\right.\right.\right.$
×$\quad\ \ \left.\left.\left.\times\cosh(\chi_{1}-\eta)\right)+(\bar{\mu}_{1}^{2}+2\bar{\mu}_{1}\bar{\mu}_{2}+\bar{\mu}_{2}+2)\right.\right.$
×$\quad\ \ \left.\left.\times\cosh(\chi_{1}-\eta)\cosh(\chi_{2}-\eta)\right)\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})\right.$
$\quad\ \ \left.+\frac{1}{2}(-\bar{\mu}_{1}+\bar{\mu}_{2})\left((\bar{\mu}_{1}+\bar{\mu}_{2})^{2}\cosh(\chi_{1}-\chi_{2})\right.\right.$
$\quad\ \ \left.\left.+(\bar{\mu}_{1}^{2}+2\bar{\mu}_{1}\bar{\mu}_{2}+\bar{\mu}_{2}+4)\cosh(2\eta+\chi_{1}+\chi_{2})\right.\right.$
@@
$\quad\ \ +\frac{-k^{\prime}\ln k^{\prime}}{\pi}\left((\bar{\mu}_{1}+\bar{\mu}_{2})\left((\bar{\mu}_{1}-\bar{\mu}_{2})^{2}\right.\right.$
×$\quad\ \ \left.\left.\times\cosh(\chi_{1}+\chi_{2})+(\bar{\mu}_{1}^{2}-2\bar{\mu}_{1}\bar{\mu}_{2}+\bar{\mu}_{2}+4)\right.\right.$
×$\quad\ \ \left.\left.\times\cosh(2\eta+\chi_{1}-\chi_{2})+4(-\bar{\mu}_{1}+\bar{\mu}_{2})\right.\right.$
×$\quad\ \ \left.\left.\times\sinh(2\eta+\chi_{1}-\chi_{2})\right)\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right.$
$\quad\ \ \left.+(\bar{\mu}_{1}+\bar{\mu}_{2})\left((\bar{\mu}_{1}-\bar{\mu}_{2})^{2}\cosh(\chi_{1}+\chi_{2})\right.\right.$
$\quad\ \ \left.\left.+(\bar{\mu}_{1}^{2}-2\bar{\mu}_{1}\bar{\mu}_{2}+\bar{\mu}_{2}+4)\cosh(-2\eta+\chi_{1}-\chi_{2})\right.\right.$
@@
×$\quad\ \ \left.\times\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right),$
$\bar{G}_{2}=\frac{-2k^{\prime}\ln k^{\prime}}{\pi}\left(\frac{1}{2}(-\bar{\mu}_{1}+\bar{\mu}_{2})\left((\bar{\mu}_{1}+\bar{\mu}_{2})\right.\right.$
×$\quad\ \ \left.\left.\times\cosh(\chi_{1}-\chi_{2})+(\bar{\mu}_{1}+\bar{\mu}_{2})\cosh(\chi_{1}+\chi_{2}-2\eta)\right.\right.$
@@
$\quad\ \ \left.+(-\bar{\mu}_{1}+\bar{\mu}_{2})\left(\sinh(\eta+\chi_{1})\cosh(\eta+\chi_{2})\right.\right.$
$\quad\ \ \left.\left.+\sinh(\eta+\chi_{2})\cosh(\eta+\chi_{1})-(\bar{\mu}_{1}+\bar{\mu}_{2})\right.\right.$
×$\quad\ \ \left.\left.\times\cosh(\eta+\chi_{1})\cosh(\eta+\chi_{2})\right)\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2})\right.$
$\quad\ \ \left.+(\bar{\mu}_{1}+\bar{\mu}_{2})\left(-\sinh(-\eta+\chi_{2})\cosh(\eta+\chi_{1})\right.\right.$
$\quad\ \ \left.\left.+\sinh(-\eta+\chi_{1})\cosh(-\eta+\chi_{2})-(\bar{\mu}_{1}-\bar{\mu}_{2})\right.\right.$
×$\quad\ \ \times\left.\left.\cosh(\eta+\chi_{1})\cosh(-\eta+\chi_{2})\right)\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right.$
$\quad\ \ \left.+(\bar{\mu}_{1}+\bar{\mu}_{2})\left(-\sinh(-\eta+\chi_{1})\cosh(\eta+\chi_{2})\right.\right.$
$\quad\ \ \left.\left.+\sinh(\eta+\chi_{2})\cosh(-\eta+\chi_{1})+(\bar{\mu}_{1}-\bar{\mu}_{2})\right.\right.$
×$\quad\ \ \times\left.\left.\cosh(-\eta+\chi_{1})\cosh(\eta+\chi_{2})\right)\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})\right),$
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (No. 12075208).
@@
are called sine-amplitude, cosine-amplitude, and delta amplitude, respectively. When the modulus $k$ tends to 00 or $1$:

### v2

@@ 6 Degenerate solutions
In this section, we discuss the degenerate cases of the periodic wave background, that is the Jacobi elliptic function with a modulus $k$ approaching either 0 or $1$, respectively. To ensure that the solutions are real after degeneracy, we consider the breather waves (53) and (65), and $\lambda\in[-k^{2}-\delta\omega_{2},+\infty)$ in both cases.
@@
$\quad\quad~\times\cosh(-\bar{\kappa}_{1}+\bar{\kappa}_{2}))^{2},$
@@
In this case, $E(k)/K(k)=0$, $F(\phi_{\gamma},k)\to\infty$ can be derived. Based on the translation invariance of the (n+1)-dimensions gKP equation (1), the half-period translation of (53) with the transformation $\eta=\eta-K(k)$ is defined to ensure the standardization of degradation solution. On account of the fact that $\gamma=F(\phi_{\gamma},k)$, for each $\lambda\in[-1-\delta\omega_{2},+\infty)$, $\chi_{1}$ can be defined according to its property HV-1969 as
@@
At the same time, employing the Poisson summation equation JPB-1982 , if $k=1$,
@@
$\quad\quad~\times\mathrm{exp}(-\bar{\kappa}_{1}))^{2},$
@@
$\quad\ \ \left.\times\mathrm{cosh(\eta-\chi_{2})}\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})+\mathrm{cosh(\eta+\chi_{1})}\right$.
$\quad\ \ \left.\times\mathrm{cosh(\eta+\chi_{2})}\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2}))+(\bar{\mu}_{1}+\bar{\mu}_{2})\right$.
$\quad\ \ \left.\times(\mathrm{cosh(\eta+\chi_{1})}\mathrm{cosh(\eta-\chi_{2})}\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right$.
@@
[…]
$\quad\ \ \left.\left.\times\cosh(\chi_{1}-\eta)\cosh(\chi_{2}-\eta)\right)\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})\right$.
[…]
@@
[…]
$\quad\ \ \left.\left.\times\sinh(2\eta+\chi_{1}-\chi_{2})\right)\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right$.
[…]
@@
$\quad\ \ \left.\times\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right),$
[…]
@@
[…]
$\quad\ \ \left.\left.\times\cosh(\eta+\chi_{1})\cosh(\eta+\chi_{2})\right)\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2})\right$.
[…]
$\quad\ \ \times\left.\left.\cosh(\eta+\chi_{1})\cosh(-\eta+\chi_{2})\right)\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})\right$.
[…]
$\quad\ \ \times\left.\left.\cosh(-\eta+\chi_{1})\cosh(\eta+\chi_{2})\right)\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})\right),$
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (No. 12475007).
Conflict of Interest
The authors declare no competing interests.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@
are called sine-amplitude, cosine-amplitude, and delta amplitude, respectively. When the modulus $k$ tends to 0 or $1$: